OpenStudy (anonymous):

Simplify the expression. -14 x 8 x 1/4 + (-22) a.-6 b.-50 c.-470 d.-138

OpenStudy (er.mohd.amir):

(-14*8*1)/4+(-22) =(-14*8)/4-22 =-28-22 =-50

OpenStudy (anonymous):

thanks can you help with more @Er.Mohd.AMIR

OpenStudy (anonymous):

Simplify. [–5 – (–3 + 4 • 5)]

OpenStudy (er.mohd.amir):

-5-(-3+20) =-5-(17) =-5+17 =12

OpenStudy (anonymous):

Simplify the expression. 3/4 [12 x 4 + (-16)] + (-20 A. 16 B. –44 C. 4 D. 44

OpenStudy (er.mohd.amir):

(3/4)(48-16) +(-20) =(3/4)(32)-20 =3*8-20 =24-20=4

OpenStudy (anonymous):

Which equations illustrate the associative property of addition? choose exactly two answeres that are correct Choose exactly two answers that are correct. A. 3/4 + 1.8 = 1.8 + 3/4 B. –3 • (5 + 8.5) = –3 • 5 + (–3) • 8.5 C. (11 + 2.2) + (–5) = 11 + [2.2 + (–5)] D. –7 + [4 + (–2.1 • 6)] = (–7 + 4) + (–2.1 • 6)

OpenStudy (er.mohd.amir):

C and D

OpenStudy (anonymous):

What number correctly completes the statement using the distributive property? 3/4 x (-5 + 4.1) = 3/4 _____ + 3/4 x 4.1

OpenStudy (er.mohd.amir):

-5

OpenStudy (er.mohd.amir):

more or finish

OpenStudy (anonymous):

more but in a sec

OpenStudy (anonymous):

Which inequality models the situation, given that d represents distance? The recommended amount of space between fence posts is greater than 6 ft. A. 6d < 0 B. d > 6 C. d < 6 D. d + 6 > 0

OpenStudy (er.mohd.amir):

B

OpenStudy (anonymous):

Which inequality models this situation? The time between two trains must be greater than 37 minutes. A. t > 37 B. 37 > t C. 37 – t = 0 D. t < 37

OpenStudy (er.mohd.amir):

A

OpenStudy (anonymous):

A frigid temperature alert is announced when the temperature is less than –10°F. Which inequality shows the temperatures when the alert will be announced? A. t > 10 B. t > –10 C. t < –10 D. t < 10

OpenStudy (er.mohd.amir):

C

OpenStudy (anonymous):

Isabella observed that everyone on her soccer team is older than 13. She wrote the inequality a > 13 to represent the possible ages of her teammates. Which numbers satisfy the inequality but make no sense in this situation? Choose all answers that are correct. A. 96 B. 13 C. 102 D. 15

OpenStudy (er.mohd.amir):

A,C,D

OpenStudy (anonymous):

I got 58%!!!

OpenStudy (er.mohd.amir):

A and C make no scene since teammates are 96 and 102 years make this

OpenStudy (er.mohd.amir):

only D make scene

OpenStudy (anonymous):

ya whatever screw you

OpenStudy (er.mohd.amir):

what

OpenStudy (er.mohd.amir):

more

OpenStudy (er.mohd.amir):

because its 2:30 am here